MICROWAVE SPECTRA AND MOLECULAR STRUCTURE OF ACETYLENEMETHYLDIOXORHENIUM, A RHENIUM METALACYCLOPROPENE^{a}

\underline{STEPHEN KUKOLICH}, BRIAN DROUIN, OLIVER INDRIS AND JENNIFER DANNEMILLER, Department of Chemistry, University of Arizona, Tucson, AZ 85721 (Present Address for B. Drouin: Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109); JOCHEN ZOLLER and WOLFGANG HERRMANN, Anorganisch-Chemisches Institut der Technischen Universität München, Lichtenbergstrasse 4, D-85747 Garching bei München, Germany.

The molecular structure for ((η^{2}-acetylene)methyldioxorhenium (ACMDO) was obtained by measuring and analysing the rotational spectra for 14 isotopomers. This appears to be the first gas-phase measurement for a rhenium metalacyclopropene, and the first structural data on this compound. This complex is closely related to reaction intermediates in methyltrioxorhenium(MTO) and OsO_{4} catalysed oxidation reactions. These reactions are very important in industrial chemical production and for syntheses of chiral products. Only a few alkyne complexes of transition metals in high oxidation states are known and even fewer have been structurally characterized. The microwave spectra were measured in the 4-11 GHz range using a Flygare-Balle-type pulsed-beam microwave spectrometer. Rotational constants and rhenium quadrupole coupling tensors were obtained for all isotopomers. The acetylene ligand structure is modified and exhibits partial sp^{2} hybridization. The C-C bond length is increased by 0.08Å to 1.29Å. The H-C-C interbond angles are reduced from 180^{∘} to 141^{∘}, and 152^{∘}. The quadrupole splitting patterns and systematic changes in the quadrupole coupling tensors were very helpful in making positive assignments.

aSupported by THE NATIONAL SCIENCE FOUNDATION